a sheet of Al foil 68.3cm L ,31.3cm W, 0.016mm Height(thick) density of 2.7g/cm 3-(squared)

determine how many moles of aluminum are present in the foil. How do you set this up ?


1
Expert's answer
2020-12-21T03:55:43-0500

L = 68.3cm

W = 31.3cm

H = 0.016mm = 0.0016 cm

Foil volume:

V(foil) = L*W*H = 68.3*31.3*0.0016 = 3.42 cm3

d - density

Foil mass:

m(foil) = V(foil)*d(foil) = 3.42*2.7 = 9.23 g

M(Al) = 27 g/mol

n(Al) = m(foil)/M(Al) = 9.23/27 = 0.342 mol

Answer: 0.342 mol